### v1

# Implementation of full and simplified likelihoods in CheckMATE Iñaki Lara and Krzysztof Rolbiecki

Abstract
We present the implementation of simplified and full likelihood models for multibin signal regions in CheckMATE. A total of 13 searches are included from ATLAS and CMS, and several methods are presented for the implementation and evaluation of likelihood functions. Statistical combinations increase the sensitivity of searches and open up the possibility of combining orthogonal search channels in the CheckMATE framework.
1 Introduction
The Run-3 at the LHC is now in full swing and by the end of the current run it is expected that the LHC would collect several hundreds of inverse femtobarns of data. Apart from the Higgs boson discovery[1, 2], which made the experimental picture of the Standard Model complete, no sign of new physics has been observed. However, the vast data has only been interpreted in the limited number of theoretical models by the LHC experiments. On many occasions, the interpretation is provided in terms of simplified models [3] which do not directly correspond to realistic TeV-scale physics. This situation has prompted several groups [4, 5, 6, 7, 8, 9] to provide computer programs that would allow reinterpretation of experimental results in terms of an arbitrary new physics model [10].
Over the years, searches for new physics were becoming increasingly sophisticated, including the full use of complicated statistical models [11] and machine learning methods [12, 13]. Instead of reporting results in a handful of bins, often optimized for a small class of models, and comparing observed and expected number of events, the experiments nowadays provide binned data in several observables in signal and control regions. On the one hand, this can significantly improve the sensitivity, but on the other hand, it greatly complicates the reinterpretation of the results. In this approach, experiments in high-energy physics test the compatibility of collision data with theoretical predictions, which can be described as a likelihood function. The function gives the probability of the data assuming a theoretical model and a certain set of parameters. The likelihood can therefore be used to constrain the parameters of the model.
In this paper, we discuss the implementation of the full and simplified likelihoods in the CheckMATE framework [14, 15, 16, 17]. For implementations in other tools, see eg. [18, 19]. In this paper, we include 9 ATLAS and 4 CMS searches. The CMS searches use the correlated background model (covariance matrix) [20]. For ATLAS searches, we use simplified [21] and full likelihood [22] frameworks. In addition, some of the implementations of the ATLAS searches include the full set of control regions. Users have several ways of controlling the actual calculation of likelihoods. The default method of evaluation is based on the package Spey [23]. We provide validation plots for each of the searches. A comparison of three methods for the calculation of limits is presented: the best-expected signal region, the full likelihood model (if available), and the simplified likelihood.
The paper is organized as follows. In Section 2, we briefly discuss statistical models and their technical implementation in CheckMATE. We provide a list of implemented searches and options available to a user to control the program’s behavior. In Section 3 we present validation plots for each of the searches discussed. A brief summary and outlook are provided in the last section.
2 Technical implementation
In this section, we introduce the methods for the implementation of simplified and full likelihoods in CheckMATE and user switches to control their execution.
2.1 ATLAS
The functionality of combining signal regions for recasting in ATLAS searches can be implemented using either the full likelihood model [22] or following a simplified approach detailed in Ref. [21]. Table 1 lists the ATLAS analyses with likelihood functionality implemented in CheckMATE. The simplified likelihood method requires background rates and uncertainties that were already available in the implemented searches. The full likelihood requires an appropriate file in the JSON format and these files were released by ATLAS for 6 searches already implemented in CheckMATE. For the searches atlas_2004_14060, atlas_2006_05880, and atlas_2111_08372 the full model files are not available but using the published data one can still perform simplified model fitting in multibin signal regions.
The full-likelihood statistical models are encoded in the JSON files by the ATLAS Collaboration. The files are not shipped with CheckMATE but are automatically downloaded during installation from the HEPData website, https://www.hepdata.net/. The information provided includes the number of background events for all signal and control regions and for each major background category separately. This results in a large number of nuisance parameters and the complexity of the procedure makes the hypothesis testing very CPU-expensive. Additionally, on the recasting side, in order to fully exploit the method, one should also implement control regions (CRs), which was not a standard approach in CheckMATE. Currently, just two searches atlas_2010_14293 and atlas_1911_06660 has a complete implementation of all CRs. In other searches, it is assumed that the contribution of signal to CRs is negligible. This assumption is not obviously fulfilled in all imaginable new physics models.
On the technical side, after the usual evaluation of events within CheckMATE, a JSON patchset is created that encapsulates signal contributions to signal regions (SRs) (and CRs if applicable). The patchset is then combined with the background-only input from ATLAS. Since the signal region names in CheckMATE do not usually exactly match the names in ATLAS conventions, an additional file is required, pyhf_conf.json, which provides a dictionary between two conventions. This is further evaluated using the package pyhf [24, 25, 26], which is a Python implementation of the HistFactory specification for binned statistical models [27, 28]. The signal strength $\mu$ is the parameter of interest, where $\mu=1$ corresponds to the nominal cross section of a tested model. Depending on the user’s choices, the output can contain information about the expected and observed upper limit on $\mu$, with 2-$\sigma$ bounds, along with the observed and expected $\mathrm{CL}_{s}$ for $\mu=1$. The default method of calculation is by using the asymptotic formulae; see [24].
By default, the above calculation will be executed using the Spey program [23] and spey-pyhf plugin [29]. Spey is a Python-based cross-platform package that allows statistical inference of hypotheses using different likelihood prescriptions.♮$\natural$1♮$\natural$1$\natural$1Installation of Spey is straightforward: pip install spey. Please refer to Spey online documentation and spey-pyhf for more details [30]. In our setup it gives a somewhat better control over the calculation than the above-mentioned CheckMATE-pyhf interface, but nevertheless the calculation is still performed in the pyhf framework. The main motivation, however, for using Spey was a possibility of combining different searches (also between experiments), which is planned in the next release of CheckMATE. In any case, a direct evaluation using pyhf and bypassing Spey also remains available.
Since the evaluation of full likelihoods is normally time-consuming, it is not practical for large scans of the parameter space. Therefore, the alternative approach to likelihood evaluation is based on the concept of simplified likelihood [21]. In this case, the background model is approximated with the total SM background rate obtained in the background-only fit in the full model. A single nuisance parameter correlated over all bins and representing post-fit background uncertainty is constrained by the unit normal distribution. The evaluation is also performed using the pyhf package.
Name Description #SR Nbin Full Ref.
[…]
Table 1: List of implemented ATLAS analyses which have likelihood-based signal regions (all searches at $\sqrt{s}=13$ TeV and $\mathcal{L}=139$ ${\text{fb}}^{-1}$).
2.2 CMS
The simplified likelihood framework was defined in Ref. [20]. This assumes correlation between background contributions that can be modeled using the multivariate Gaussian distribution:
$\mathcal{L}_{S}(\mu,\bm{\theta})=\prod_{i=1}^{N}\frac{(\mu\cdot s_{i}+b_{i}+\theta_{i})^{n_{i}}e^{-(\mu\cdot s_{i}+b_{i}+\theta_{i})}}{n_{i}!}\cdot\exp\left(-\frac{1}{2}\bm{\theta}^{T}\mathbf{V}^{-1}\bm{\theta}\right)$ (1)
where the product runs over all bins and $\mu$ is the signal strength (and the Parameter of Interest - POI), $n_{i}$ the observed number of events, $s_{i}$ an expected number of signal events, $b_{i}$ an expected number of background events, $\theta_{i}$ a background nuisance parameter and $\mathbf{V}$ the covariance matrix. It is implemented using the covariance matrices provided by the CMS Collaboration which are included in the CheckMATE distribution in the JSON format. Evaluation of the above model is performed using the Spey package and the default_pdf.correlated_background method.
Name Description Nbin Ref.
cms_1908_04722 Search for supersymmetry in final states with jets and $E_{\mathrm{T}}^{\mathrm{miss}}$ 174 [40]
cms_1909_03460 Search for supersymmetry with $M_{\mathrm{T2}}$ variable in final states with jets and $E_{\mathrm{T}}^{\mathrm{miss}}$ 282 [41]
cms_2107_13021 Search for new particles in events with energetic jets and large $E_{\mathrm{T}}^{\mathrm{miss}}$ 66 [42]
cms_2205_09597 Search for production of charginos and neutralinos in final states containing hadronic decays of $WW$, $WZ$, or $WH$ and $E_{\mathrm{T}}^{\mathrm{miss}}$ 35 [42]
Table 2: List of implemented CMS analyses which have likelihood-based signal regions (all searches at $\sqrt{s}=13$ TeV and $\mathcal{L}=139$ ${\text{fb}}^{-1}$).
2.3 CheckMATE parameters
CheckMATE provides several switches and parameters to control details of statistical evaluation. These are summarized in Tab. 3. The switches are divided into two groups: one providing a control of what statistical tests are performed and the other to control different modes of calculation. By default no statistical evaluation is performed. For the sake of speed and stability one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information. Generally the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures. By choosing a select switch users can control which statistics are calculated. If no explicit choice is made the observed upper limit will be calculated. Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming. The option -so can be used to request calculation of statistics for previous CheckMATE runs (it requires presence of the evaluation/total_results.dat file in the output directory).
The second group of parameters is used to choose a method of calculation of requested statistics for the ATLAS searches (it does not affect the calculation for the CMS searches as described in the previous Section). For the default method, simple, calculation is performed using simplified likelihood and the CheckMATE-pyhf interface. The full switch chooses a calculation using the full likelihood and the CheckMATE-Spey interface. Finally, the fullpyhf switch requests calculation sing the full likelihood and CheckMATE-pyhf interface (this is somewhat less flexible regarding the output compared to the previous options). In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
Parameter card Terminal X Description and available choices
Multibin: X -mb X none No signal region combination is performed (default).
select Calculates user selected statistics.
scan Calculates observed $\mathrm{CL}_{s}$; fast and reliable for quick assessment of exclusion.
detailed Calculates observed and expected upper limits and $\mathrm{CL}_{s}$.
Expected: False -exp Selects calculation of expected limits.
CLs: False -mbcls Selects calculation of $\mathrm{CL}_{s}$.
Uplim: False -uplim Selects calculation of upper limits.
Statonly: False -so Calculates statistical combinations without event-level analysis provided the analysis and evaluation steps were already completed.
Model: X -mod X simple The simplified likelihood model for ATLAS searches (default).
full The Spey interface to the full likelihood model for ATLAS searches.
fullpyhf The full likelihood model for ATLAS searches with pyhf interface.
Backend: X numpy Different backends for calculations in the full model.
pytorch Pytorch backend (if available)
tensorflow Tensorflow backend (if available)
jax Jax backend (if available)
Table 3: Summary of options related to multibin signal regions.
The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file. In order to follow the progress of calculation the observed limits are also displayed on screen for each of the signal regions. The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
$\mu<1\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Exclued}$
$\mu\geq 1\ \Longrightarrow\ \texttt{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}Allowed}.$
If the results for upper limit are not available the decision is made using the observed $\mathrm{CL}_{s}$ statistics at the 95% confidence level:
$\mathrm{CL}_{s}<0.05\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Exclued}$
$\mathrm{CL}_{s}\geq 0.05\ \Longrightarrow\ \texttt{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}Allowed}.$
2.4 Evaluation time
Four different backends for calculation of full likelihood models in the pyhf/spey setup are supported: numpy [43], tensorflow [44], pytorch [45], and jax [46]. The numpy backend does not require additional system components to be installed but it generally is not recommended due to long computation time. When no backend is explicitly selected, CheckMATE will check if the jax backend is available. The jax backend was also successfully tested with NVIDIA CUDA library and GPU support.
An example calculation was performed for the evaluation of the full likelihood model in the ATLAS search atlas_2101_01629, which is a moderately complex one of those implemented. The tests were performed on three Ubuntu 22.04.5 systems: A - Intel i7-4770 CPU 3.4 GHz, 4/8 cores/threads;♮$\natural$2♮$\natural$2$\natural$2It is really old. B - Intel i7-8565U CPU 1.8 GHz, 4/8 cores/threads and NVIDIA GeForce MX150; C - Intel i7-7700 CPU 3.4 GHz, 4/8 cores/threads. Jax, pytorch and tensorflow take advantage of the multithread computation even when no GPUs are present. On system B the CUDA 12.8 toolkit was installed for the GPU computation. In Table 4 we show the results. Note that the system B was tested in two configurations, with and without, GPUs. The best performance is obtained with jax and pytorch. For comparison, we also provide evaluation times for the $\mathrm{CL}_{s}$ observed limit performed on system A.
As already mentioned, for large scans of parameter space it may not be practical to perform full likelihood evaluation. The simplified models improve computation and the detailed calculation of all signal strength limits and $\mathrm{CL}_{s}$ takes in our example 16 seconds for the jax backend and 64 seconds for the numpy backend. Finally, if the calculation of $\mathrm{CL}_{s}$ is sufficient, it only takes about 1 second.
upper limits observed and expected $\mathrm{CL}_{s}$
Backend CPU A CPU B GPU B CPU C CPU A
numpy 11092 12621 N/A 8879 636
jax 723 526 353 311 55
pytorch 583 805 N/A 404 57
tensorflow 2426 2920 1906 1808 156
Table 4: Performance comparison (time in seconds) of different pyhf backends for calculation of the full likelihood in the analysis atlas_2101_01629. Systems A and C run on (multithread) CPU, while system B was tested on CPU and GPU. As a reference in the last column the computation time of $\mathrm{CL}_{s}$ observed limit is shown.
3 Validation
Multibin signal regions are currently available in 9 ATLAS and 4 CMS analyses, as listed in Tabs. 1 and 2. The searches are based on the full Run 2 luminosity of about 140 ${\text{fb}}^{-1}$ at the center-of-mass energy $\sqrt{s}=13\ \mathrm{TeV}$. In this section we briefly introduce each of the searches and provide validation examples. When possible we compare full vs. simplified likelihood approach, and provide examples how the multibin approach improves exclusions compared to the best-signal-region approach. The best-SR (BSR) is defined as the signal region with the best expected sensitivity.
The validation procedure is organized as follows. For SUSY processes events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state. The NNPDF23LO [50, 51, 52] parton distribution function (PDF) set is used. The events are then interfaced to Pythia-8.2 [53, 54] for modeling of decays, hadronization and showering. The matrix element and parton shower matching was done using the CKKW-L [55] prescription and a matching scale of 1/4 of the SUSY particle mass was set. Inclusive signal cross sections for production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at next-to-next-to-leading-logarithm (approximate NNLO+NNLL) [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], following recommendations of Ref. [67]. The signal cross sections for production of sleptons, charginos and neutralinos are computed at next-to-leading order plus next-to-leading-log precision using Resummino [68, 69, 70, 71, 72, 73, 74]. This setup generally follows procedures employed within LHC experiments to obtain exclusion limits used in the validation process.
For the purpose of validation, we present comparisons between CheckMATE and official ATLAS result of both expected and observed exclusion limits. Where available, the results for simplified and full likelihood are provided. In addition, we also include exclusion contours using the best expected signal region. Note that this usually is based on comparing observed signal with the post-fit background-only prediction in a given bin, which typically has reduced uncertainty, and hence can be quite an aggressive approach. Also, a limit obtained with the full likelihood model is not a combination of such bin-by-bin measurements and cannot be compared in a straightforward way. On the other hand, the simplified likelihood will be such a combination in most cases and it is not surprising that it will produce stronger constraints than the full likelihood in some searches, see also Ref. [23] The latter generally shows weaker sensitivity than multibin histogram shape-fits, but can be advantageous in terms of evaluation speed.
We follow the same key for the plots throughout this section. The black solid (dashed) line is for ATLAS/CMS observed (expected) limit. On some occasions this is accompanied by theoretical (experimental) ±$\pm 1\sigma$ uncertainty. The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for observed and expected limits, respectively. Finally, the blue and green lines denote limits obtained using the simplified and full likelihoods, respectively.
3.1 atlas_1908_03122 (SUSY-2018-31)
This is a search [31] for bottom squark production in final states containing Higgs bosons, $b$-jets, and missing transverse momentum. The Higgs boson is reconstructed from two $b$-tagged jets. The final states contain at least 3 (SRC) or 4 (SRA, SRB) $b$-jets, no leptons, and large missing transverse momentum. The signal region A is divided into 3 bins according to effective mass, $m_{\textrm{eff}}$, and the signal region C is divided into 4 bins of missing transverse energy significance, $\mathcal{S}$. Thus, both SRA and SRC allow for a shape-fit analysis. The full likelihood model is provided.
The validation plots for this search are presented in Fig. 1. We compare bottom squark pair production, $pp\to\tilde{b}_{1}\tilde{b}_{1}^{*}$, and two distinct mass spectra. The first assumes that the mass of the lightest supersymmetric particle (LSP) is $m_{\tilde{\chi}_{1}^{0}}=60$ GeV and the other assumes the mass difference between neutralinos $\Delta m(\tilde{\chi}_{2}^{0},\tilde{\chi}_{1}^{0})=130$ GeV. Generally a good agreement between CheckMATE and ATLAS is observed for the shape-fit method with CheckMATE being slightly weaker. Notably, there is a very good agreement between simplified and full models. The best SR method performs well for the model with fixed LSP mass, however, in the second scenario its exclusion strength can be up to 3 times weaker than shape-fits.
Figure 1: Validation plots for the search atlas_1908_03122 (SUSY-2018-31). Top row: the model with $m_{\tilde{\chi}_{1}^{0}}=60$ GeV; bottom row: the model with $\Delta m(\tilde{\chi}_{2}^{0},\tilde{\chi}_{1}^{0})=130$ GeV. Left panels: observed limits; right panels: expected limits.
3.2 atlas_1908_08215 (SUSY-2018-32)
This is a search [32] for production of electroweakinos and sleptons. The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered. Events are first separated into SF and DF categories and further subdivided by the multiplicity of the non-b-tagged jets. A combined multi-bin SR is defined out of the 36 exclusive binned signal regions. The full likelihood statistical model is provided.
The validation plots for the production of slepton pairs are shown in Fig. 2. Both the full- and simplified likelihood models agree very well with the ATLAS result. The BSR method gives visibly weaker constraints.
Figure 2: Validation plots for the search atlas_1908_08215 (SUSY-2018-33), slepton pair production. Left panels: observed limits; right panels: expected limits.
3.3 atlas_1911_06660 (SUSY-2018-04)
This is a search [33] for production of staus in final states with two hadronic $\tau$-leptons and missing transverse momentum. Two orthogonal signal regions can be combined in a fit for which the full likelihood model is provided. The CheckMATE implementation includes one of the multijet control regions (CR-A), for which a significant contribution is expected, up to 30%, from signal events.
The validation plots in Fig. 3 show a similar sensitivity of the best signal region, as well as full and simplified likelihoods. In each case, good agreement with the ATLAS result is observed. The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably inside $1\sigma$ experimental uncertainty. It should be noted that including the control region in the full likelihood fit was essential in obtaining the acceptable agreement.
Figure 3: Validation plots for the search atlas_1911_06660 (SUSY-2018-04). Left panels: observed limits; right panels: expected limits.
3.4 atlas_1911_12606 (SUSY-2018-16)
This is a search [34] for production of electroweakinos and sleptons in scenarios with compressed mass spectra. The final states contain two low $p_{T}$ leptons (opposite sign and same or different flavor). Sensitivity of the search relies on additional initial state radiation (ISR) jets which give transverse boost to the final state particles and adds missing transverse momentum. There are two multibin signal regions implemented in CheckMATE:♮$\natural$3♮$\natural$3$\natural$3The vector boson fusion (VBF) SRs are currently not available. SR-EWK targeting production of electroweakinos and divided into 44 bins according to the lepton pair invariant mass, $E_{\mathrm{T}}^{\mathrm{miss}}$, and lepton flavor; SR-S targeting production of sleptons and divided into 32 bins. The full likelihood model is provided.
In Figure 4, we show validation plots for slepton search regions. The electron and muon channels are shown separately. Very good agreement between ATLAS and CheckMATE is obtained for the shape-fit analysis. Additionally, by comparing red exclusion curves obtained using the best-SR-method, we find a clear benefit of using multibin signal regions. Since there is a negligible difference between expected and observed results (both for ATLAS and CheckMATE) we only show the observed limits.
In Figure 5, we show validation plots for the wino-like chargino-neutralino pair production (EW search region). The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row. Electron and muon channels are shown separately. These examples show a danger of using a simplified likelihood model, which clearly produces too strong constraints.
Figure 4: Validation plots (observed limits) for the search atlas_1911_12606 (SUSY-2018-16), slepton pair production. Left panel: smuon search; right panel selectron search.
Figure 5: Validation plots for the search atlas_1911_12606 (SUSY-2018-16). Top row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$; bottom row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$. Left panels: observed limits; right panels: expected limits.
3.5 atlas_2004_14060 (SUSY-2018-12)
This is a search [35] for hadronically decaying supersymmetric partners of top quark, and up-type, 3rd generation scalar leptoquark. The final states consist of several jets, 0 leptons, and large $E_{\mathrm{T}}^{\mathrm{miss}}$. The requirements for $b$ jets vary between signal regions. There are 3 multibin signal regions: SRA-B which targets scenarios with highly boosted top quarks in the final state and is divided into 6 bins according to an invariant mass of a large-$R$ jet; SRC for scenarios with 3-body decays of stops which is divided into 5 bins according to a recursive jigsaw reconstruction technique variable $R_{\textrm{ISR}}$ [75]; SRD for scenarios with compressed spectra and 4-body decays of stops which is divided into 3 bins according to the number of identified $b$-jets. No likelihood model is provided, and only simplified likelihood fitting is available.
The validation plots for this search are shown in Figure 6. We include the full range of stop masses and decay modes. A good agreement across the parameter plane is observed. For the high-stop-mass region the CheckMATE result is slightly weaker than ATLAS, clearly extending the exclusion reach compared to single SR exclusion. In the more compressed scenarios, the simplified fit gives a slightly too strong exclusion in some regions of the parameter space, however, this is below 30% difference in the exclusion strength.
Figure 6: Validation plots for the search atlas_2004_14060 (SUSY-2018-12). Left panels: observed limits; right panels: expected limits.
3.6 atlas_2006_05880 (SUSY-2018-21)
This is a search [36] for supersymmetric partners of top quark decaying to a Higgs or $Z$ boson. The final state Higgs boson is reconstructed from a pair of $b$-jets while the $Z$ boson from a same-flavor opposite-sign dilepton pair. There are 3 multibin signal regions that are shape fits in $E_{\mathrm{T}}^{\mathrm{miss}}$, $E_{\mathrm{T}}^{\mathrm{miss}}$-significance and $p_{\mathrm{T}}$ of the $Z$ candidate. No likelihood model is provided, and only simplified likelihood model is available.
The validation plots in Figure 7 again show a good agreement, with CheckMATE somewhat weaker than ATLAS in the high stop-mass region. The observed limits clearly extend the reach compared to the best SR, increasing sensitivity by up to factor 2.
Figure 7: Validation plots for the search atlas_2006_05880 (SUSY-2018-31). Left panels: observed limits; right panels: expected limits.
3.7 atlas_2010_14293 (SUSY-2018-22)
This is a search [37] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons and missing transverse momentum. There are three multibin signal regions in this search: MB-SSd which targets production of squarks and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets production of gluinos and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets. The control regions are also implemented what gives an opportunity to fully exploit the full likelihood model provided by the ATLAS collaboration.
Figure 8 shows validation plots for the squark pair-production process. The simplified shape-fit has excellent agreement with the official result, while the full likelihood model gives a somewhat weaker exclusion. In either case there is a clearly visible advantage over the best-SR method.
Figure 8: Validation plots for the search atlas_2010_14293 (SUSY-2018-22), squark pair production - 8-fold degeneracy. Left panels: observed limits; right panels: expected limits.
3.8 atlas_2101_01629 (SUSY-2018-10)
This is a search [38] for squarks and gluinos in final states with one isolated lepton, jets, and missing transverse momentum. Benchmark models assume long decay chains for squarks and gluinos with charginos, neutralinos and gauge bosons in intermediate states that give rise to the final state lepton. There is one multibin signal region that combines 26 bins defined according to the number of jets, number of identified $b$-jets, and $m_{\textrm{eff}}$. The full likelihood model is provided.
The validation plots for gluino pair production, followed by the decay to intermediate mass chargino are shown in Fig. 9. The shape of the ATLAS exclusion line is well reproduced for the likelihood models, whereas the BSR exclusion is significantly weaker.
Figure 9: Validation plots for the search atlas_2101_01629, gluino pair production and the decay chain: $\tilde{g}\to q\bar{q}^{\prime}\tilde{\chi}_{1}^{\pm}\to q\bar{q}^{\prime}W\tilde{\chi}_{1}^{0}$ with $x=(m(\tilde{\chi}_{1}^{\pm})-m(\tilde{\chi}_{1}^{0}))/(m(\tilde{g})-m(\tilde{\chi}_{1}^{0}))=1/2$.
3.9 atlas_2111_08372 (HIGG-2018-26)
This is a search [39] for invisible decays of a Higgs boson or dark matter (DM) particles, $\chi$, produced in association with a $Z$ boson. The final state $Z$ boson is reconstructed from a same-flavor opposite-sign dilepton pair. One multibin signal region, optimized for the 2HDM+$a$ model [76], is implemented in CheckMATE. It is divided into 22 bins according to transverse mass, $m_{\mathrm{T}}$. The implementation is based on the simplified likelihood model.
As can be seen in Figure 10 we find excellent agreement between the result of the simplified shape-fit and the official result. The best-SR exclusion is significantly weaker. It should be noted that in the parts of the plot with large $m_{a}$ and/or $m_{A}$ the exclusion is driven by the shape-fit in the range of large $m_{T}$, which increases the sensitivity by factor $\sim 2$.
Figure 10: Validation plots for the search atlas_2111_08372 (HIGG-2018-26) in the 2HDM+$a$ model, $pp\to Z(\to\ell^{+}\ell^{-})a(\to\chi\chi)$. Left panel: observed limits; right panel: expected limits.
3.10 cms_1908_04722 (SUS-19-006)
This is a search [40] for supersymmetric particles in final states with jets ($\geq 2$) and missing transverse momentum. The main discriminating variables are $H_{T}$ and $H_{\mathrm{T}}^{\mathrm{miss}}$, the scalar sum of jet transverse momenta, and the magnitude of the vector sum of jet transverse momenta, respectively. Only the signal regions for prompt production are currently implemented in CheckMATE. The search combines 174 bins in the simplified likelihood framework with the covariance matrix provided by the CMS collaboration. Individual bins are defined according to the number of jets and $b$-jets, $H_{\mathrm{T}}$, and $H_{\mathrm{T}}^{\mathrm{miss}}$. In addition, there are 12 aggregate signal regions defined.
In Figure 11 we provide validation plots for two cases: squark pair production (single generation) and gluino pair production (followed by $\tilde{g}\to t\bar{t}\tilde{\chi}_{1}^{0}$ decay). In both cases, we observe good agreement, within 1-sigma uncertainty) between the CheckMATE likelihood model and the CMS results. In case of gluinos there is a significant improvement with respect to the best-SR result, especially in the large gluino-mass limit. It is worth noting the small production cross section in this region: $\sigma\sim 0.6$ fb at $m_{\tilde{g}}\sim 2.2$ TeV.
Figure 11: Validation plots for the search cms_1908_04722, squark pair production (1-fold degeneracy; upper row), gluino pair production with $\tilde{g}\to t\bar{t}\tilde{\chi}^{0}_{1}$ (lower row). The dotted lines around ATLAS limit denote 1-sigma uncertainty: theoretical for the observed limit and experimental for the expected limit. Left panels: observed limits; right panels: expected limits.
3.11 cms_1909_03460 (SUS-19-005)
This is a search [40] for supersymmetric particles (squarks and gluinos) in final states with jets ($\geq 2$) and a large missing transverse momentum. The events are required to have at least one energetic jet. For the inclusive search, events are required to have at least 2 jets and signal regions are based on the number of jets, $b$-jets, $H_{T}$, and the $M_{T2}$ variable calculated using two pseudojets. Only the signal regions for prompt production are currently implemented in CheckMATE. There are 21 inclusive aggregate signal regions and 282 exclusive bins which are combined into two multi-bin signal regions using a covariance matrix. Due to numerical problems with the entire matrix $282\times 282$, we split the calculation into two multibin signal regions (corresponding to final states with low- and high-$H_{T}$, respectively) that are marginally correlated [77]. However, these two signal regions can be further combined by a user with an option SRCombination.
In Figure 12 we show the validation plot for gluino pair production followed by decay $\tilde{g}\to t\bar{t}\tilde{\chi}_{1}^{0}$. The result is based on the above-mentioned combination of multibin signal regions. There is good agreement between CheckMATE (green lines) and ATLAS (black lines). There is also a clear advantage of the multibin fir over the best-SR (here the aggregate SRs were used).
Figure 12: Validation plots for the search cms_1909_03460. The dotted lines around ATLAS limit denote 1-sigma uncertainty: theoretical for the observed limit and experimental for the expected limit. Left panel: observed limits; right panel: expected limits.
3.12 cms_2107_13021 (EXO-20-004)
This is a search [42] for new particles in the final states with at least one jet, no leptons, and missing transverse momentum. A main focus of the analysis are invisible particles that can be dark-matter candidates and that are produced with at least one ISR jet. The simplified likelihood fit is performed on 66 bins: 3 sets for different data-taking periods and divided according to missing transverse energy.
For validation, we compare exclusion limits for the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane. The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively. The validation plot, Figure 13, shows again good agreement between the CheckMATE and CMS limits.
Figure 13: Validation plots for the search cms_2107_13021 and the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane. The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively. The diagonal line indicates $m_{\textrm{med}}=2m_{\textrm{DM}}$.
3.13 cms_2205_09597 (SUS-21-002)
This is a search [78] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, $WZ$ and $WH$. This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets. The individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates and missing transverse momentum. The simplified likelihood is built out of 35 signal regions. Additionally, there are 4 aggregate signal regions defined.
Currently, the implementation is validated for the production of charginos with $W$ bosons in the final state. As can be seen in Fig. 14 there is good agreement between CheckMATE and the CMS exclusion contours. The multibin fit clearly improves sensitivity, the red vs. green line in the left panel of Fig. 14. The remaining signal regions for $Z$ and $h$ channels are implemented, however, due to missing mistag rates for different types of final states, they should be used with caution.
Figure 14: Validation plots for the search cms_2205_09597 and chargino pair production followed by $\tilde{\chi}^{\pm}_{1}\to W^{\pm}\tilde{\chi}_{1}^{0}$.
4 Conclusions and outlook
In this paper, we presented a version of CheckMATE, which significantly updates the database of available experimental searches. All included searches are based on the full data set of LHC Run 2. Here, we describe the implementation of 9 ATLAS searches and 4 CMS searches. Another update comes with the inclusion of statistical models, which in a number of examples significantly improve sensitivity and expand exclusion limits. This implementation includes full and simplified likelihoods provided by the ATLAS collaboration, and simplified correlated background models provided by CMS. With a number of options available, it allows users to fully control the numerical evaluation. CheckMATE can be downloaded from the GitHub repository:
https://github.com/CheckMATE2/checkmate2/  .
More information, including previous versions, expanded validation notes can be found at:
http://checkmate.hepforge.org/  .
In future releases, we plan to further expand CheckMATE’s capabilities by allowing statistical combination of different searches and orthogonal signal regions. In a separate note we will cover another new development to CheckMATE, which is an interface to machine learning methods which are becoming increasingly common tool in experimental analyses at the LHC.
Acknowledgements
The authors thank Zirui Wang for his help and additional input for Ref. [39]. This work was partially supported by the National Science Centre, Poland under grant 2019/35/B/ST2/02008. The authors have also received funding from the Norwegian Financial Mechanism 2014-2021, grant 2019/34/H/ST2/00707.

### v2

@@
 The CMS searches use the correlated background model (covariance matrix) [20].
-For ATLAS searches, we use simplified [21] and full likelihood [22] frameworks.
+For ATLAS searches, we use the simplified [21] and the full likelihood [22] frameworks.
 In addition, some of the implementations of the ATLAS searches include the full set of control regions.
@@
 The default method of calculation is by using the asymptotic formulae; see [24].
-By default, the above calculation will be executed using the Spey program [23] and spey-pyhf plugin [29].
-Spey is a Python-based cross-platform package that allows statistical inference of hypotheses using different likelihood prescriptions.♮$\natural$1♮$\natural$1$\natural$1Installation of Spey is straightforward: pip install spey.
-Please refer to Spey online documentation and spey-pyhf for more details [30].
-In our setup it gives a somewhat better control over the calculation than the above-mentioned CheckMATE-pyhf interface, but nevertheless the calculation is still performed in the pyhf framework.
-The main motivation, however, for using Spey was a possibility of combining different searches (also between experiments), which is planned in the next release of CheckMATE.
-In any case, a direct evaluation using pyhf and bypassing Spey also remains available.
+[…]
 Since the evaluation of full likelihoods is normally time-consuming, it is not practical for large scans of the parameter space.
@@
 […]
-[…]
+atlas_2111_08372 Search for associated production of a Z boson with an invisibly decaying Higgs boson or dark matter candidates $2 \usym1$F5F4 [39]
 Table 1: List of implemented ATLAS analyses which have likelihood-based signal regions (all searches at $\sqrt{s}=13$ TeV and $\mathcal{L}=139$ ${\text{fb}}^{-1}$).
@@
 cms_2107_13021 Search for new particles in events with energetic jets and large $E_{\mathrm{T}}^{\mathrm{miss}}$ 66 [42]
-cms_2205_09597 Search for production of charginos and neutralinos in final states containing hadronic decays of $WW$, $WZ$, or $WH$ and $E_{\mathrm{T}}^{\mathrm{miss}}$ 35 [42]
+cms_2205_09597 Search for production of charginos and neutralinos in final states containing hadronic decays of $WW$, WZWZ, or $WH$ and $E_{\mathrm{T}}^{\mathrm{miss}}$ 35 [42]
 Table 2: List of implemented CMS analyses which have likelihood-based signal regions (all searches at $\sqrt{s}=13$ TeV and $\mathcal{L}=139$ ${\text{fb}}^{-1}$).
@@
 The switches are divided into two groups: one providing a control of what statistical tests are performed and the other to control different modes of calculation.
-By default no statistical evaluation is performed.
-For the sake of speed and stability one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information.
-Generally the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures.
+By default, no statistical evaluation is performed.
+For the sake of speed and stability, one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information.
+Generally, the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures.
 By choosing a select switch users can control which statistics are calculated.
-If no explicit choice is made the observed upper limit will be calculated.
+If no explicit choice is made, the observed upper limit will be calculated.
 Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming.
@@
 The full switch chooses a calculation using the full likelihood and the CheckMATE-Spey interface.
-Finally, the fullpyhf switch requests calculation sing the full likelihood and CheckMATE-pyhf interface (this is somewhat less flexible regarding the output compared to the previous options).
+Finally, the fullpyhf switch requests the calculation using the full likelihood and CheckMATE-Pyhf interface (this is somewhat less flexible regarding the output compared to the previous options).
 In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
@@
 fullpyhf The full likelihood model for ATLAS searches with Pyhf interface.
-Backend: X numpy Different backends for calculations in the full model.
+Backend: X numpy Different backends for calculations in the full model; Numpy is always available.
 pytorch Pytorch backend (if available)
@@
 The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file.
-In order to follow the progress of calculation the observed limits are also displayed on screen for each of the signal regions.
+In order to follow the progress of the calculation, the observed limits are also displayed on screen for each of the signal regions.
 The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
@@
 The Jax backend was also successfully tested with NVIDIA CUDA library and GPU support.
-An example calculation was performed for the evaluation of the full likelihood model in the ATLAS search atlas_2101_01629, which is a moderately complex one of those implemented.
-The tests were performed on three Ubuntu 22.04.5 systems: A - Intel i7-4770 CPU 3.4 GHz, 4/8 cores/threads;♮$\natural$2♮$\natural$2$\natural$2It is really old.
-B - Intel i7-8565U CPU 1.8 GHz, 4/8 cores/threads and NVIDIA GeForce MX150; C - Intel i7-7700 CPU 3.4 GHz, 4/8 cores/threads.
-Jax, pytorch and tensorflow take advantage of the multithread computation even when no GPUs are present.
-On system B the CUDA 12.8 toolkit was installed for the GPU computation.
-In Table 4 we show the results.
-Note that the system B was tested in two configurations, with and without, GPUs.
-The best performance is obtained with jax and pytorch.
-For comparison, we also provide evaluation times for the $\mathrm{CL}_{s}$ observed limit performed on system A.
+[…]
 As already mentioned, for large scans of parameter space it may not be practical to perform full likelihood evaluation.
@@
 The validation procedure is organized as follows.
-For SUSY processes events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state.
+For SUSY processes, events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state.
 The NNPDF23LO [50, 51, 52] parton distribution function (PDF) set is used.
-The events are then interfaced to Pythia-8.2 [53, 54] for modeling of decays, hadronization and showering.
-The matrix element and parton shower matching was done using the CKKW-L [55] prescription and a matching scale of 1/4 of the SUSY particle mass was set.
+The events are then interfaced to Pythia 8.3 [53, 54] for modeling of decays, hadronization and showering.
+The matrix element and parton shower matching for ATLAS searches was performed using the CKKW-L [55] prescription and a matching scale of 1/4 of the SUSY particle mass was set.
+For CMS searches, the MLM [56, 48] matching scheme was used.
 Inclusive signal cross sections for production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at next-to-next-to-leading-logarithm (approximate NNLO+NNLL) [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], following recommendations of Ref.
@@
 This setup generally follows procedures employed within LHC experiments to obtain exclusion limits used in the validation process.
+Events for the simplified scenario of DM production in the CMS search EXO-20-004 [42] were generated at NLO with MadGraph5_aMC@NLO 3.1.0 and the DMsimp model [76].
+Up to 2 additional partons were included and the FxFx [77] jet matching scheme was used.
+Samples for the 2HDM+a model and validation of HIGG-2018-26 [39] were simulated internally in CheckMATE interfaced to MadGraph5_aMC@NLO 3.4.1 and hadronized with Pythia 8.2.
 For the purpose of validation, we present comparisons between CheckMATE and official ATLAS result of both expected and observed exclusion limits.
@@
 The black solid (dashed) line is for ATLAS/CMS observed (expected) limit.
-On some occasions this is accompanied by theoretical (experimental) ±$\pm 1\sigma$ uncertainty.
+On some occasions, this is accompanied by theoretical (experimental) $\pm 1\sigma$ uncertainty.
 The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for observed and expected limits, respectively.
-Finally, the blue and green lines denote limits obtained using the simplified and full likelihoods, respectively.
+Finally, the blue and green lines denote the limits obtained using the simplified and full likelihoods, respectively.
 3.1 atlas_1908_03122 (SUSY-2018-31)
@@
 The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered.
-Events are first separated into SF and DF categories and further subdivided by the multiplicity of the non-b-tagged jets.
+Events are first separated into SF and DF categories and are further subdivided by the multiplicity of the non-b-tagged jets.
 A combined multi-bin SR is defined out of the 36 exclusive binned signal regions.
@@
 In each case, good agreement with the ATLAS result is observed.
-The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably inside $1\sigma$ experimental uncertainty.
+The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably within the experimental $1\sigma$ uncertainty.
 It should be noted that including the control region in the full likelihood fit was essential in obtaining the acceptable agreement.
@@
 3.4 atlas_1911_12606 (SUSY-2018-16)
-This is a search [34] for production of electroweakinos and sleptons in scenarios with compressed mass spectra.
-The final states contain two low $p_{T}$ leptons (opposite sign and same or different flavor).
-Sensitivity of the search relies on additional initial state radiation (ISR) jets which give transverse boost to the final state particles and adds missing transverse momentum.
-There are two multibin signal regions implemented in CheckMATE:♮$\natural$3♮$\natural$3$\natural$3The vector boson fusion (VBF) SRs are currently not available.
-SR-EWK targeting production of electroweakinos and divided into 44 bins according to the lepton pair invariant mass, $E_{\mathrm{T}}^{\mathrm{miss}}$, and lepton flavor; SR-S targeting production of sleptons and divided into 32 bins.
-The full likelihood model is provided.
+[…]
 In Figure 4, we show validation plots for slepton search regions.
@@
 In Figure 5, we show validation plots for the wino-like chargino-neutralino pair production (EW search region).
-The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row.
-Electron and muon channels are shown separately.
+The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row; see eg.
+Ref.
+[78] for a discussion of CP properties of the neutralino sector and consequences for the decay kinematics.
+The electron and muon channels are shown separately.
 These examples show a danger of using a simplified likelihood model, which clearly produces too strong constraints.
@@
 There are 3 multibin signal regions that are shape fits in $E_{\mathrm{T}}^{\mathrm{miss}}$, $E_{\mathrm{T}}^{\mathrm{miss}}$-significance and $p_{\mathrm{T}}$ of the $Z$ candidate.
-No likelihood model is provided, and only simplified likelihood model is available.
+No likelihood model is provided, and only a simplified likelihood model is available.
 The validation plots in Figure 7 again show a good agreement, with CheckMATE somewhat weaker than ATLAS in the high stop-mass region.
@@
 There are three multibin signal regions in this search: MB-SSd which targets production of squarks and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets production of gluinos and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets.
-The control regions are also implemented what gives an opportunity to fully exploit the full likelihood model provided by the ATLAS collaboration.
-Figure 8 shows validation plots for the squark pair-production process.
+The control regions are also implemented, which gives an opportunity to fully exploit the full likelihood model provided by the ATLAS Collaboration.
+Figure 8 shows the validation plots for the squark pair-production process.
 The simplified shape-fit has excellent agreement with the official result, while the full likelihood model gives a somewhat weaker exclusion.
-In either case there is a clearly visible advantage over the best-SR method.
+In either case, there is a clearly visible advantage over the best-SR method.
 Figure 8: Validation plots for the search atlas_2010_14293 (SUSY-2018-22), squark pair production - 8-fold degeneracy. Left panels: observed limits; right panels: expected limits.
@@
 The implementation is based on the simplified likelihood model.
-As can be seen in Figure 10 we find excellent agreement between the result of the simplified shape-fit and the official result.
+As can be seen in Figure 10 we find excellent agreement between the simplified shape fit result and the official result.
 The best-SR exclusion is significantly weaker.
-It should be noted that in the parts of the plot with large $m_{a}$ and/or $m_{A}$ the exclusion is driven by the shape-fit in the range of large $m_{T}$, which increases the sensitivity by factor $\sim 2$.
+It should be noted that in the parts of the plot with large $m_{a}$ and/or $m_{A}$ the exclusion is driven by the shape fit in the range of large $m_{T}$, which increases the sensitivity by factor $\sim 2$.
 Figure 10: Validation plots for the search atlas_2111_08372 (HIGG-2018-26) in the 2HDM+$a$ model, $pp\to Z(\to\ell^{+}\ell^{-})a(\to\chi\chi)$.
@@
 The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively.
-The validation plot, Figure 13, shows again good agreement between the CheckMATE and CMS limits.
+The validation plot, Figure 13, again shows good agreement between the CheckMATE and CMS limits.
 Figure 13: Validation plots for the search cms_2107_13021 and the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane.
@@
 3.13 cms_2205_09597 (SUS-21-002)
-This is a search [78] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, $WZ$ and $WH$.
+This is a search [82] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, WZWZ and $WH$.
 This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets.

### v3

@@
-Implementation of full and simplified likelihoods in CheckMATE Iñaki Lara and Krzysztof Rolbiecki
+ul. Pasteura 5, PL-02-093 Warsaw, Poland
 Abstract
@@
 1 Introduction
-The Run-3 at the LHC is now in full swing and by the end of the current run it is expected that the LHC would collect several hundreds of inverse femtobarns of data.
+The Run 3 at the Large Hadron Collider (LHC) is now in full swing, and by the end of the current run it is expected that the LHC would collect several hundreds of inverse femtobarns of data.
 Apart from the Higgs boson discovery[2, 52], which made the experimental picture of the Standard Model complete, no sign of new physics has been observed.
@@
 Over the years, searches for new physics were becoming increasingly sophisticated, including the full use of complicated statistical models [59] and machine learning methods [15, 25].
-Instead of reporting results in a handful of bins, often optimized for a small class of models, and comparing observed and expected number of events, the experiments nowadays provide binned data in several observables in signal and control regions.
+Instead of reporting results in a handful of bins, often optimized for a small class of models, and comparing observed and expected numbers of events, the experiments nowadays provide binned data in several observables in signal and control regions.
 On the one hand, this can significantly improve the sensitivity, but on the other hand, it greatly complicates the reinterpretation of the results.
@@
 In this paper, we discuss the implementation of the full and simplified likelihoods in the CheckMATE framework [65, 77, 63, 64].
-For implementations in other tools, see eg.
+For implementations in other tools, see, e.g.
 [18, 16].
@@
 The CMS searches use the correlated background model (covariance matrix) [55].
-For ATLAS searches, we use the simplified [21] and the full likelihood [22] frameworks.
+For ATLAS searches, we use the simplified [74] and full likelihood [88] frameworks.
 In addition, some of the implementations of the ATLAS searches include the full set of control regions.
 Users have several ways of controlling the actual calculation of likelihoods.
-The default method of evaluation is based on the package Spey [23].
+The default method of evaluation is based on the Spey package [26].
 We provide validation plots for each of the searches.
@@
 The full likelihood requires an appropriate file in the JSON format and these files were released by ATLAS for 6 searches already implemented in CheckMATE.
-For the searches atlas_2004_14060, atlas_2006_05880, and atlas_2111_08372 the full model files are not available but using the published data one can still perform simplified model fitting in multibin signal regions.
+For the searches atlas_2004_14060, atlas_2006_05880, and atlas_2111_08372 the full model files are not available but using the published data one can still perform a simplified model fitting in multibin signal regions.
 The full-likelihood statistical models are encoded in the JSON files by the ATLAS Collaboration.
-The files are not shipped with CheckMATE but are automatically downloaded during installation from the HEPData website, https://www.hepdata.net/.
+The files are not shipped with CheckMATE but are automatically downloaded during installation from the HEPData website [48, 83], https://www.hepdata.net/.
 The information provided includes the number of background events for all signal and control regions and for each major background category separately.
 This results in a large number of nuisance parameters and the complexity of the procedure makes the hypothesis testing very CPU-expensive.
-Additionally, on the recasting side, in order to fully exploit the method, one should also implement control regions (CRs), which was not a standard approach in CheckMATE.
-Currently, just two searches atlas_2010_14293 and atlas_1911_06660 has a complete implementation of all CRs.
+Additionally, on the recasting side, in order to fully exploit the method, one should also implement control regions (CR), which was not a standard approach in CheckMATE.
+Currently, only two searches atlas_2010_14293 and atlas_1911_06660 have complete implementation of all CRs.
 In other searches, it is assumed that the contribution of signal to CRs is negligible.
@@
 On the technical side, after the usual evaluation of events within CheckMATE, a JSON patchset is created that encapsulates signal contributions to signal regions (SRs) (and CRs if applicable).
-The patchset is then combined with the background-only input from ATLAS.
+The patchset is then combined with the background-only model from ATLAS.
 Since the signal region names in CheckMATE do not usually exactly match the names in ATLAS conventions, an additional file is required, pyhf_conf.json, which provides a dictionary between two conventions.
-This is further evaluated using the package Pyhf [24, 25, 26], which is a Python implementation of the HistFactory specification for binned statistical models [27, 28].
+The likelihood is further evaluated using the Pyhf [57, 72, 73] package, which is a Python implementation of the HistFactory specification for binned statistical models [58, 28].
 The signal strength $\mu$ is the parameter of interest, where $\mu=1$ corresponds to the nominal cross section of a tested model.
-Depending on the user’s choices, the output can contain information about the expected and observed upper limit on $\mu$, with 2-$\sigma$ bounds, along with the observed and expected $\mathrm{CL}_{s}$ for $\mu=1$.
-The default method of calculation is by using the asymptotic formulae; see [24].
-[…]
-Since the evaluation of full likelihoods is normally time-consuming, it is not practical for large scans of the parameter space.
-Therefore, the alternative approach to likelihood evaluation is based on the concept of simplified likelihood [21].
-In this case, the background model is approximated with the total SM background rate obtained in the background-only fit in the full model.
-A single nuisance parameter correlated over all bins and representing post-fit background uncertainty is constrained by the unit normal distribution.
-The evaluation is also performed using the Pyhf package.
-Name Description #SR Nbin Full Ref.
+Depending on the user’s choices, the output can contain information about the expected and observed upper limit on $\mu$, with 2-$\sigma$ bounds, along with the observed and expected $\mathrm{CL}_{s}$ for $\mu=1$, where
+$\mathrm{CL}_{s}=\frac{\mathrm{CL}_{s+b}}{\mathrm{CL}_{b}}$ (1)
+and $\mathrm{CL}_{s+b}$, $\mathrm{CL}_{b}$ are the $p$-values of the signal and the null hypotheses, respectively. Asymptotic formulae are the default calculation method; see [57].
 […]
 […]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-atlas_2111_08372 Search for associated production of a Z boson with an invisibly decaying Higgs boson or dark matter candidates $2 \usym1$F5F4 [39]
+Name Description #SR Nbin Full Ref.
+atlas_1908_03122 Search for bottom squarks in final states with Higgs bosons, $b$-jets and $E_{\mathrm{T}}^{\mathrm{miss}}$ 2 7
+[3]
+atlas_1908_08215 Search for electroweak production of charginos and sleptons in final states with 2 leptons and $E_{\mathrm{T}}^{\mathrm{miss}}$ 1 52
+[6]
+atlas_1911_06660 Search for direct stau production in events with two hadronic taus 1 2
+[5]
+atlas_1911_12606 Search for electroweak production of supersymmetric particles with compressed mass spectra 2 76
+[8]
+atlas_2004_14060 Search for stops in hadronic final states with $E_{\mathrm{T}}^{\mathrm{miss}}$ 3 14
+[4]
+atlas_2006_05880 Search for top squarks in events with a Higgs or $Z$ boson 3 23
+[7]
+atlas_2010_14293 Search for squarks and gluinos in final states with jets and $E_{\mathrm{T}}^{\mathrm{miss}}$ 3 60
+[9]
+atlas_2101_01629 Search for squarks and gluinos in final states with one isolated lepton, jets, and $E_{\mathrm{T}}^{\mathrm{miss}}$ 1 26
+[10]
+atlas_2111_08372 Search for associated production of a $Z$ boson with an invisibly decaying Higgs boson or dark matter candidates 1 22
+[11]
 Table 1: List of implemented ATLAS analyses which have likelihood-based signal regions (all searches at $\sqrt{s}=13$ TeV and $\mathcal{L}=139$ ${\text{fb}}^{-1}$).
@@
 $\mathcal{L}_{S}(\mu,\bm{\theta})=\prod_{i=1}^{N}\frac{(\mu\cdot s_{i}+b_{i}+\theta_{i})^{n_{i}}e^{-(\mu\cdot s_{i}+b_{i}+\theta_{i})}}{n_{i}!}\cdot\exp\left(-\frac{1}{2}\bm{\theta}^{T}\mathbf{V}^{-1}\bm{\theta}\right)$ (2)
-where the product runs over all bins and $\mu$ is the signal strength (and the Parameter of Interest - POI), $n_{i}$ the observed number of events, $s_{i}$ an expected number of signal events, $b_{i}$ an expected number of background events, $\theta_{i}$ a background nuisance parameter and $\mathbf{V}$ the covariance matrix.
-It is implemented using the covariance matrices provided by the CMS Collaboration which are included in the CheckMATE distribution in the JSON format.
-Evaluation of the above model is performed using the Spey package and the default_pdf.correlated_background method.
+where the product runs over all bins and $\mu$ is the signal strength (and the Parameter of Interest - POI), $n_{i}$ the observed number of events, $s_{i}$ the expected number of signal events, $b_{i}$ the expected number of background events, $\theta_{i}$ a background nuisance parameter, and $\mathbf{V}$ the covariance matrix.
+It is implemented using the covariance matrices provided by the CMS Collaboration, which are included in the CheckMATE distribution in the JSON format.
+The evaluation of the above model is performed using the Spey package and the default_pdf.correlated_background method.
 Name Description Nbin Ref.
@@
 2.3 CheckMATE parameters
-CheckMATE provides several switches and parameters to control details of statistical evaluation.
+CheckMATE provides several switches and parameters to control the details of the statistical evaluation.
 These are summarized in Tab.
 3.
-The switches are divided into two groups: one providing a control of what statistical tests are performed and the other to control different modes of calculation.
+The switches are divided into two groups: one providing a control of what statistical tests are performed, and the other to control different modes of calculation.
 By default, no statistical evaluation is performed.
-For the sake of speed and stability, one switch, scan, provides a quick and reliable way of obtaining Allowed/Excluded result but with limited additional information.
-Generally, the available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures.
-By choosing a select switch users can control which statistics are calculated.
+For the sake of speed and stability, one switch, scan, provides a quick and reliable way of obtaining an Allowed/Excluded result but with limited additional information.
+Generally, available statistics include $\mathrm{CL}_{s}$ tests and calculation of upper limits on signal strength, both of which can be obtained as observed and/or expected measures.
+By choosing a select switch, users can control which statistics are calculated.
 If no explicit choice is made, the observed upper limit will be calculated.
 Finally, the detailed switch can be used to request calculation of all available statistics, but it should be noted that its execution can be time consuming.
-The option -so can be used to request calculation of statistics for previous CheckMATE runs (it requires presence of the evaluation/total_results.dat file in the output directory).
+The option -so can be used to request the calculation of statistics for previous CheckMATE runs (it requires the presence of the evaluation/total_results.dat file in the output directory).
 The second group of parameters is used to choose a method of calculation of requested statistics for the ATLAS searches (it does not affect the calculation for the CMS searches as described in the previous Section).
-For the default method, simple, calculation is performed using simplified likelihood and the CheckMATE-Pyhf interface.
-The full switch chooses a calculation using the full likelihood and the CheckMATE-Spey interface.
-Finally, the fullpyhf switch requests the calculation using the full likelihood and CheckMATE-Pyhf interface (this is somewhat less flexible regarding the output compared to the previous options).
+For the default method, the full switch chooses the calculation using the full likelihood and the CheckMATE-Spey interface.
+With the simple switch, the calculation is performed using simplified likelihood and the CheckMATE-Pyhf interface.
+Finally, the fullpyhf switch requests the calculation using the full likelihood and CheckMATE-Pyhf interface (this is somewhat less flexible with regard to the output compared to the previous options).
 In any case, users should remember that the full likelihood calculation can be time consuming if many searches and signal regions are requested.
@@
 Statonly: False -so Calculates statistical combinations without event-level analysis provided the analysis and evaluation steps were already completed.
-Model: X -mod X simple The simplified likelihood model for ATLAS searches (default).
-full The Speyinterface to the full likelihood model for ATLAS searches.
+full The Speyinterface to the full likelihood model for ATLAS searches (default).
+Model: X -mod X simple The simplified likelihood model for ATLAS searches.
 fullpyhf The full likelihood model for ATLAS searches with Pyhf interface.
@@
 The results of the calculation for each of the multibin signal regions and all requested analyses are stored in the multibin_limits/results.dat file.
-In order to follow the progress of the calculation, the observed limits are also displayed on screen for each of the signal regions.
+In order to follow the progress of the calculation, the observed limits are also displayed on the screen for each of the signal regions.
 The final evaluation is decided using the upper limit on the signal strength $\mu$ (if calculated):
-$\mu<1\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Exclued}$
+$\mu<1\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Excluded}$
 $\mu\geq 1\ \Longrightarrow\ \texttt{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}Allowed}.$
-If the results for upper limit are not available the decision is made using the observed $\mathrm{CL}_{s}$ statistics at the 95% confidence level:
-$\mathrm{CL}_{s}<0.05\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Exclued}$
+If the results for the upper limit are not available, the decision is made using the observed $\mathrm{CL}_{s}$ statistics at the 95% confidence level:
+$\mathrm{CL}_{s}<0.05\ \Longrightarrow\ \texttt{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Excluded}$
 $\mathrm{CL}_{s}\geq 0.05\ \Longrightarrow\ \texttt{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}Allowed}.$
 2.4 Evaluation time
-Four different backends for calculation of full likelihood models in the pyhf/Spey setup are supported: Numpy [43], Tensorflow [44], Pytorch [45], and Jax [46].
-The Numpy backend does not require additional system components to be installed but it generally is not recommended due to long computation time.
+Four different backends for the calculation of full likelihood models in the pyhf/Spey setup are supported: Numpy [86], Tensorflow [96], Pytorch [87], and Jax [76].
+The Numpy backend does not require additional system components to be installed, but it generally is not recommended due to long computation time.
 When no backend is explicitly selected, CheckMATE will check if the Jax backend is available.
@@
 As already mentioned, for large scans of parameter space it may not be practical to perform full likelihood evaluation.
-The simplified models improve computation and the detailed calculation of all signal strength limits and $\mathrm{CL}_{s}$ takes in our example 16 seconds for the Jax backend and 64 seconds for the Numpy backend.
+The simplified models improve the computation and detailed calculation of all signal strength limits and $\mathrm{CL}_{s}$ takes in our example 16 seconds for the Jax backend and 64 seconds for the Numpy backend.
 Finally, if the calculation of $\mathrm{CL}_{s}$ is sufficient, it only takes about 1 second.
@@
 The searches are based on the full Run 2 luminosity of about 140 ${\text{fb}}^{-1}$ at the center-of-mass energy $\sqrt{s}=13\ \mathrm{TeV}$.
-In this section we briefly introduce each of the searches and provide validation examples.
-When possible we compare full vs.
-simplified likelihood approach, and provide examples how the multibin approach improves exclusions compared to the best-signal-region approach.
+In this section, we briefly introduce each of the searches and provide validation examples.
+When possible, we compare the full and simplified likelihood approaches, and provide examples of how the multibin evaluation improves exclusion limits compared to the best-signal-region approach.
 The best-SR (BSR) is defined as the signal region with the best expected sensitivity.
 The validation procedure is organized as follows.
-For SUSY processes, events are generated using MadGraph5_aMC@NLO 3.1.0 [47, 48, 49] with up to two additional partons in the final state.
+For SUSY processes, events are generated using MadGraph5_aMC@NLO 3.1.0 [21, 23, 22] and the MSSM_SLHA2 [66] UFO [60] model with up to two additional partons in the final state.
 The NNPDF23LO [31, 47, 32] parton distribution function (PDF) set is used.
-The events are then interfaced to Pythia 8.3 [53, 54] for modeling of decays, hadronization and showering.
+The events are then interfaced with Pythia 8.3 [91, 44] to model decays, hadronization, and showering.
 The matrix element and parton shower matching for ATLAS searches was performed using the CKKW-L [82] prescription and a matching scale of 1/4 of the SUSY particle mass was set.
 For CMS searches, the MLM [84, 23] matching scheme was used.
-Inclusive signal cross sections for production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at next-to-next-to-leading-logarithm (approximate NNLO+NNLL) [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67], following recommendations of Ref.
+In the final step, a fast detector simulation is performed using Delphes 3.5 [61].
+Jets are reconstructed using FastJet [51] and the anti-$k_{t}$ algorithm [50].
+Inclusive signal cross sections for the production of squarks and gluinos are obtained at the approximate next-to-next-to-leading order with soft gluon resummation at the next-to-next-to-leading-logarithm order (approximate NNLO+NNLL) [38, 34, 80, 79, 39, 41, 42, 37, 35, 40, 36], following the recommendations of Ref.
 [49].
@@
 Up to 2 additional partons were included and the FxFx [69] jet matching scheme was used.
-Samples for the 2HDM+a model and validation of HIGG-2018-26 [39] were simulated internally in CheckMATE interfaced to MadGraph5_aMC@NLO 3.4.1 and hadronized with Pythia 8.2.
-For the purpose of validation, we present comparisons between CheckMATE and official ATLAS result of both expected and observed exclusion limits.
+The samples for the 2HDM+a model [33] and the validation of HIGG-2018-26 [11] were simulated internally in CheckMATE interfaced with MadGraph5_aMC@NLO 3.4.1 and hadronized with Pythia 8.2.
+The sampling of parameter space generally follows grids provided by experiments in the supplementary materials on HEPData.
+The number of Monte-Carlo (MC) events per point varies; however, close to exclusion lines at mass limits, we aim at samples 10-20 times larger than the nominal luminosity derived values.
+The exclusion lines are then interpolated based on the grids of simulated parameter points, with additional sampling points occasionally added.
+MC errors are consistently included in the pyhf patch files.
+For the purpose of validation, we present comparisons between CheckMATE and the official ATLAS result of both expected and observed exclusion limits.
 Where available, the results for simplified and full likelihood are provided.
-In addition, we also include exclusion contours using the best expected signal region.
+In addition, we also include exclusion contours using the best-expected signal region.
 Note that this usually is based on comparing observed signal with the post-fit background-only prediction in a given bin, which typically has reduced uncertainty, and hence can be quite an aggressive approach.
 Also, a limit obtained with the full likelihood model is not a combination of such bin-by-bin measurements and cannot be compared in a straightforward way.
-On the other hand, the simplified likelihood will be such a combination in most cases and it is not surprising that it will produce stronger constraints than the full likelihood in some searches, see also Ref.
+On the other hand, the simplified likelihood will be such a combination in most cases and it is not surprising that it will produce stronger constraints than the full likelihood in some searches; see also Ref.
 [26] The latter generally shows weaker sensitivity than multibin histogram shape-fits, but can be advantageous in terms of evaluation speed.
 We follow the same key for the plots throughout this section.
-The black solid (dashed) line is for ATLAS/CMS observed (expected) limit.
+The black solid (dashed) line is for the observed (expected) ATLAS/CMS limit.
 On some occasions, this is accompanied by theoretical (experimental) $\pm 1\sigma$ uncertainty.
-The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for observed and expected limits, respectively.
+The red lines are used for the exclusion obtained with CheckMATE using the best signal region method, solid and dashed for the observed and expected limits, respectively.
 Finally, the blue and green lines denote the limits obtained using the simplified and full likelihoods, respectively.
@@
 Thus, both SRA and SRC allow for a shape-fit analysis.
-The full likelihood model is provided.
+Three full likelihood models are provided for the signal regions SRA, SRB, and SRC.
 The validation plots for this search are presented in Fig.
@@
 Generally a good agreement between CheckMATE and ATLAS is observed for the shape-fit method with CheckMATE being slightly weaker.
-Notably, there is a very good agreement between simplified and full models.
+In particular, there is a very good agreement between simplified and full models.
 The best SR method performs well for the model with fixed LSP mass, however, in the second scenario its exclusion strength can be up to 3 times weaker than shape-fits.
@@
 3.2 atlas_1908_08215 (SUSY-2018-32)
-This is a search [32] for production of electroweakinos and sleptons.
+This is a search [6] for the production of electroweakinos and sleptons.
 The final states with two leptons (opposite sign and same or different flavor SF/DF) and $E_{\mathrm{T}}^{\mathrm{miss}}$are considered.
@@
 The BSR method gives visibly weaker constraints.
+Additional validation material for CheckMATE, including cut-flows, can be found in [95].
 Figure 2: Validation plots for the search atlas_1908_08215 (SUSY-2018-33), slepton pair production. Left panels: observed limits; right panels: expected limits.
 3.3 atlas_1911_06660 (SUSY-2018-04)
-This is a search [33] for production of staus in final states with two hadronic $\tau$-leptons and missing transverse momentum.
+This is a search [5] for the production of staus in final states with two hadronic $\tau$-leptons and missing transverse momentum.
 Two orthogonal signal regions can be combined in a fit for which the full likelihood model is provided.
-The CheckMATE implementation includes one of the multijet control regions (CR-A), for which a significant contribution is expected, up to 30%, from signal events.
+The implementation includes one of the multijet control regions (CR-A), for which a significant contribution is expected, up to 30%, from signal events.
 The validation plots in Fig.
@@
 The observed exclusion limits slightly exceed the ATLAS limit, whereas for the expected limits, the best agreement is obtained for the full-likelihood approach, and comfortably within the experimental $1\sigma$ uncertainty.
-It should be noted that including the control region in the full likelihood fit was essential in obtaining the acceptable agreement.
-Figure 3: Validation plots for the search atlas_1911_06660 (SUSY-2018-04). Left panels: observed limits; right panels: expected limits.
+A comparison of observed 95% CL upper limits on cross section for several representative parameter points is shown in Tab.
+5.
+In the bottom plot of Fig.
+3 we additionally compare the expected limits obtained with and without the inclusion of control regions.
+Note that for this search, the signal contribution to the control regions is not negligible.
+Without control regions, the search over-constrains the parameter space, although the discrepancy remains within the 1-$\sigma$ expected uncertainty.
+Figure 3: Validation plots for the search atlas_1911_06660 (SUSY-2018-04).
+Left panels: observed limits; right panels: expected limits.
+In the bottom row: a comparison between the observed limits derived with and without control regions.
+$(m_{\tilde{\tau}_{1}},m_{\tilde{\chi}_{1}^{0}})$ [GeV] full simple ATLAS
+$(320,120)$ $3.7$ $4.0$ $4.6$
+$(320,160)$ $5.2$ $5.8$ $5.0$
+$(400,80)$ $1.9$ $1.6$ $2.0$
+$(440,80)$ $1.5$ $1.4$ $1.7$
+Table 5: Comparison of observed 95% CL upper limits on cross section in fb for stau pair production in the search atlas_1911_06660 (SUSY-2018-04); full and simple columns correspond to the limits obtained with CheckMATE using the full and simplified likelihood, respectively.
 3.4 atlas_1911_12606 (SUSY-2018-16)
 […]
-In Figure 4, we show validation plots for slepton search regions.
+In Figure 4, we show validation plots for slepton search signal regions.
 The electron and muon channels are shown separately.
@@
 Additionally, by comparing red exclusion curves obtained using the best-SR-method, we find a clear benefit of using multibin signal regions.
-Since there is a negligible difference between expected and observed results (both for ATLAS and CheckMATE) we only show the observed limits.
-In Figure 5, we show validation plots for the wino-like chargino-neutralino pair production (EW search region).
-The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row; see eg.
+Since there is a negligible difference between expected and observed results (both for ATLAS and CheckMATE), we only show the observed limits.
+In Figure 5, we show validation plots for the production of wino-like chargino-neutralino pairs (EW search region).
+The searches for different parities of LSP and NLSP are shown separately: upper row $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$ (same parity) and $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$ (opposite parity) in the lower row; see, e.g.
 Ref.
 [53] for a discussion of CP properties of the neutralino sector and consequences for the decay kinematics.
-The electron and muon channels are shown separately.
-These examples show a danger of using a simplified likelihood model, which clearly produces too strong constraints.
-Figure 4: Validation plots (observed limits) for the search atlas_1911_12606 (SUSY-2018-16), slepton pair production. Left panel: smuon search; right panel selectron search.
+Clearly, the simplified likelihood has a problem with reproducing the complicated shape of the observed limit due to significant excesses and deficits in several signal regions.
+Note that a similar conclusion was obtained by ATLAS in Ref.
+[74].
+Therefore, for this search, we advise against using the simplified likelihood in routine runs.
+The agreement between the full likelihood and the ATLAS observed limit for the case of same parity neutralinos (top row) is not ideal as well, even though the expected limits comparison is rather encouraging and within 1-$\sigma$ uncertainty, see the bottom plot in Fig.
+5.
+The search is also included in HackAnalysis, cf.
+Ref.
+[14], SModelS [19], and MadAnalysis [24].
+Additional validation material for CheckMATE, including cut-flows, can be found in [93].
+Figure 4: Validation plots (left: observed limits; right: expected limits) for the search atlas_1911_12606 (SUSY-2018-16), slepton pair production.
+Top row: smuon search; bottom row: selectron search.
+Left panel: observed limits; right panel: expected limits.
 Figure 5: Validation plots for the search atlas_1911_12606 (SUSY-2018-16).
-Top row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$; bottom row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$.
+Top row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})>0$; middle row: $m(\tilde{\chi}_{2}^{0})\times m(\tilde{\chi}_{1}^{0})<0$.
 Left panels: observed limits; right panels: expected limits.
+In the bottom row comparison of expected 1-$\sigma$ bands of ATLAS and CheckMATE full likelihood model for the same-parity case.
 3.5 atlas_2004_14060 (SUSY-2018-12)
@@
 The requirements for $b$ jets vary between signal regions.
-There are 3 multibin signal regions: SRA-B which targets scenarios with highly boosted top quarks in the final state and is divided into 6 bins according to an invariant mass of a large-$R$ jet; SRC for scenarios with 3-body decays of stops which is divided into 5 bins according to a recursive jigsaw reconstruction technique variable $R_{\textrm{ISR}}$ [79]; SRD for scenarios with compressed spectra and 4-body decays of stops which is divided into 3 bins according to the number of identified $b$-jets.
+There are 3 multibin signal regions: SRA-B that targets scenarios with highly boosted top quarks in the final state and is divided into 6 bins according to an invariant mass of a large-$R$ jet; SRC, for scenarios with 3-body decays of stops, that is divided into 5 bins according to a recursive jigsaw reconstruction technique variable $R_{\textrm{ISR}}$ [75]; SRD for scenarios with compressed spectra and 4-body decays of stops which is divided into 3 bins according to the number of identified $b$-jets.
 No likelihood model is provided, and only simplified likelihood fitting is available.
+The combinations are performed separately for each signal category (corresponding to three likelihood models), A-B, C, and D, as described in the ATLAS paper.
+This avoids overlaps and accidental correlations.
 The validation plots for this search are shown in Figure 6.
 We include the full range of stop masses and decay modes.
-A good agreement across the parameter plane is observed.
+Good agreement across the parameter plane is observed.
 For the high-stop-mass region the CheckMATE result is slightly weaker than ATLAS, clearly extending the exclusion reach compared to single SR exclusion.
-In the more compressed scenarios, the simplified fit gives a slightly too strong exclusion in some regions of the parameter space, however, this is below 30% difference in the exclusion strength.
+In more compressed scenarios, the simplified fit gives a slightly too strong exclusion in some regions of the parameter space; however, this is below 30% difference in the exclusion strength.
+Additional validation material can be found in [92].
 Figure 6: Validation plots for the search atlas_2004_14060 (SUSY-2018-12). Left panels: observed limits; right panels: expected limits.
 3.6 atlas_2006_05880 (SUSY-2018-21)
-This is a search [36] for supersymmetric partners of top quark decaying to a Higgs or $Z$ boson.
-The final state Higgs boson is reconstructed from a pair of $b$-jets while the $Z$ boson from a same-flavor opposite-sign dilepton pair.
-There are 3 multibin signal regions that are shape fits in $E_{\mathrm{T}}^{\mathrm{miss}}$, $E_{\mathrm{T}}^{\mathrm{miss}}$-significance and $p_{\mathrm{T}}$ of the $Z$ candidate.
-No likelihood model is provided, and only a simplified likelihood model is available.
+[…]
 The validation plots in Figure 7 again show a good agreement, with CheckMATE somewhat weaker than ATLAS in the high stop-mass region.
 The observed limits clearly extend the reach compared to the best SR, increasing sensitivity by up to factor 2.
-Figure 7: Validation plots for the search atlas_2006_05880 (SUSY-2018-31). Left panels: observed limits; right panels: expected limits.
+Figure 7: Validation plots for the search atlas_2006_05880 (SUSY-2018-21). Left panels: observed limits; right panels: expected limits.
 3.7 atlas_2010_14293 (SUSY-2018-22)
-This is a search [37] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons and missing transverse momentum.
-There are three multibin signal regions in this search: MB-SSd which targets production of squarks and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets production of gluinos and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets.
+This is a search [9] for squarks (1st and 2nd generation) and gluinos in final states with 2-6 jets, 0 leptons, and missing transverse momentum.
+There are three multibin signal regions in this search: MB-SSd which targets squark pair production and is divided into 24 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets; MB-GGd which targets gluino production and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$; MB-C which targets compressed spectra and is divided into 18 bins according to $m_{\textrm{eff}}$, $E_{\mathrm{T}}^{\mathrm{miss}}/\sqrt{H_{\mathrm{T}}}$ and the number of jets.
 The control regions are also implemented, which gives an opportunity to fully exploit the full likelihood model provided by the ATLAS Collaboration.
@@
 In either case, there is a clearly visible advantage over the best-SR method.
+In Appendix we provide detailed signal yields for the multibin SRs of this search.
+Additional validation material (cut flows) can be found in [94].
 Figure 8: Validation plots for the search atlas_2010_14293 (SUSY-2018-22), squark pair production - 8-fold degeneracy. Left panels: observed limits; right panels: expected limits.
 3.8 atlas_2101_01629 (SUSY-2018-10)
-This is a search [38] for squarks and gluinos in final states with one isolated lepton, jets, and missing transverse momentum.
+This is a search [10] for squarks and gluinos in final states with an isolated lepton, jets, and missing transverse momentum.
 Benchmark models assume long decay chains for squarks and gluinos with charginos, neutralinos and gauge bosons in intermediate states that give rise to the final state lepton.
-There is one multibin signal region that combines 26 bins defined according to the number of jets, number of identified $b$-jets, and $m_{\textrm{eff}}$.
+There is a multibin signal region that combines 26 bins defined according to the number of jets, the number of identified $b$-jets, and $m_{\textrm{eff}}$.
 The full likelihood model is provided.
-The validation plots for gluino pair production, followed by the decay to intermediate mass chargino are shown in Fig.
+The validation plots for gluino pair production, followed by the decay to intermediate mass chargino, are shown in Fig.
 9.
@@
 The final state $Z$ boson is reconstructed from a same-flavor opposite-sign dilepton pair.
-One multibin signal region, optimized for the 2HDM+$a$ model [80], is implemented in CheckMATE.
-It is divided into 22 bins according to transverse mass, $m_{\mathrm{T}}$.
+A multibin signal region, optimized for the 2HDM+$a$ model [33], is implemented in CheckMATE.
+It is divided into 22 bins according to the transverse mass, $m_{\mathrm{T}}$.
 The implementation is based on the simplified likelihood model.
@@
 Only the signal regions for prompt production are currently implemented in CheckMATE.
-There are 21 inclusive aggregate signal regions and 282 exclusive bins which are combined into two multi-bin signal regions using a covariance matrix.
-Due to numerical problems with the entire matrix $282\times 282$, we split the calculation into two multibin signal regions (corresponding to final states with low- and high-$H_{T}$, respectively) that are marginally correlated [81].
-However, these two signal regions can be further combined by a user with an option SRCombination.
+There are 21 inclusive aggregate signal regions and 282 exclusive bins, which are combined into two multi-bin signal regions using a covariance matrix.
+Due to numerical problems with the entire matrix $282\times 282$, we split the calculation into two multibin signal regions (corresponding to final states with low and high-$H_{T}$, respectively) that are marginally correlated [85].
+However, these two signal regions can be further combined by a user with the SRCombination input parameter, and the UnCorrStatisticsCombiner function of the Spey package.
 In Figure 12 we show the validation plot for gluino pair production followed by decay $\tilde{g}\to t\bar{t}\tilde{\chi}_{1}^{0}$.
@@
 For validation, we compare exclusion limits for the vector mediator model in the $m_{\textrm{mediator}}$$m_{\textrm{DM}}$ plane.
-The mediator-quarks and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively.
+The mediator-quark and mediator-DM couplings are fixed to $g_{q}=0.25$ and $g_{\chi}=1$, respectively.
 The validation plot, Figure 13, again shows good agreement between the CheckMATE and CMS limits.
@@
 3.13 cms_2205_09597 (SUS-21-002)
-This is a search [82] for charginos and neutralinos, that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, WZWZ and $WH$.
+This is a search [98] for charginos and neutralinos that looks for final states with large missing transverse momentum and pairs of hadronically decaying bosons $WW$, WZWZ and $WH$.
 This search makes use of specific algorithms (taggers) defined to identify $W$ boson/Higgs boson candidates out of the identified large-radius signal jets.
-The individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates and missing transverse momentum.
+Individual bins are defined orthogonally based on the number of $b$-tagged jets, the number of identified $W$, $Z$, and Higgs boson candidates, and missing transverse momentum.
 The simplified likelihood is built out of 35 signal regions.
@@
 4 Conclusions and outlook
-In this paper, we presented a version of CheckMATE, which significantly updates the database of available experimental searches.
+In this paper, we present a version of CheckMATE, which significantly updates the database of available experimental searches.
 All included searches are based on the full data set of LHC Run 2.
@@
 Another update comes with the inclusion of statistical models, which in a number of examples significantly improve sensitivity and expand exclusion limits.
-This implementation includes full and simplified likelihoods provided by the ATLAS collaboration, and simplified correlated background models provided by CMS.
+This implementation includes the full and simplified likelihoods provided by the ATLAS Collaboration and the simplified correlated background models provided by CMS.
 With a number of options available, it allows users to fully control the numerical evaluation.
@@
 https://github.com/CheckMATE2/checkmate2/ .
-More information, including previous versions, expanded validation notes can be found at:
+More information, including previous versions and expanded validation notes, can be found at:
 http://checkmate.hepforge.org/ .
-In future releases, we plan to further expand CheckMATE’s capabilities by allowing statistical combination of different searches and orthogonal signal regions.
-In a separate note we will cover another new development to CheckMATE, which is an interface to machine learning methods which are becoming increasingly common tool in experimental analyses at the LHC.
+In future releases, we plan to further expand the CheckMATEcapabilities by allowing a statistical combination of different searches and orthogonal signal regions.
+In a separate note, we will cover another new development to CheckMATE, which is an interface to machine learning methods which are becoming increasingly common tools in experimental analyses at the LHC.
 Acknowledgements
@@
 [11].
-This work was partially supported by the National Science Centre, Poland under grant 2019/35/B/ST2/02008.
+This work was supported by the National Science Centre, Poland under grant 2019/35/B/ST2/02008 and OpenMAPP project under CHIST-ERA programme (grant No.
+NCN 2022/04/Y/ST2/00186).
 The authors have also received funding from the Norwegian Financial Mechanism 2014-2021, grant 2019/34/H/ST2/00707.
+This version of the article has been accepted for publication after peer review, but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections.
+The Version of Record is available online at: https://doi.org/10.1140/epjc/s10052-026-15414-8.
+Appendix
+Event yields for atlas_2010_14293 (SUSY-2018-22)
+In Tables 68 we provide a detailed comparison of the event yields in the signal regions of the atlas_2010_14293 (SUSY-2018-22) search where the MC sample size in CheckMATE was 100k.
+A good agreement (within statistical uncertainties) between ATLAS and CheckMATE is observed except for the 4-jet bins in the squark search region MB-SSd.
+However, the test sample, the squark pair production with direct decays $\tilde{q}\to q\tilde{\chi}_{1}^{0}$, at the lowest parton order does not contribute to these bins and requires 2 additional jets from ISR/FSR.
+It is therefore very sensitive to details in MC modeling and is difficult to reproduce in recasts.
+These discrepancies do not affect the fit for this signal process due to the small yields compared to expected backgrounds.
+Bin CheckMATE ATLAS ATLAS/CheckMATE
+MB-SSd-2-1000-10 $0.94\pm 0.033$ $\pm 0.15 1.1$$1\pm 0.16$
+MB-SSd-2-1000-16 $0.74\pm 0.029$ $0.59\pm 0.1$ $0.8\pm 0.14$
+MB-SSd-2-1000-22 $0.41\pm 0.022$ $0.27\pm 0.071$ $0.66\pm 0.17$
+MB-SSd-2-1600-10 $3.8\pm 0.066$ $3.7\pm 0.26$ $0.96\pm 0.069$
+MB-SSd-2-1600-16 $2.2\pm 0.05$ $2\pm 0.18$ $0.91\pm 0.085$
+MB-SSd-2-1600-22 $4\pm 0.067$ $3.5\pm 0.26$ $0.88\pm 0.064$
+MB-SSd-2-2200-16 $1.8\pm 0.045$ $2\pm 0.2$ $1.1\pm 0.11$
+MB-SSd-2-2200-22 $7.8\pm 0.094$ $8.4\pm 0.39$ $1.1\pm 0.051$
+MB-SSd-2-2800-16 $1.4\pm 0.04$ $1.3\pm 0.16$ $0.95\pm 0.11$
+MB-SSd-2-2800-22 $5.9\pm 0.082$ $6\pm 0.33$ $1\pm 0.055$
+MB-SSd-2-3400-22 $1.2\pm 0.038$ $0.93\pm 0.14$ $0.77\pm 0.12$
+MB-SSd-2-3400-28 $6\pm 0.055 2.2$$2\pm 0.19$ $0.84\pm 0.075$
+MB-SSd-2-4000-22 $0.47\pm 0.024$ $0.22\pm 0.063$ $0.47\pm 0.13$
+MB-SSd-2-4000-28 $0.89\pm 0.033$ $0.67\pm 0.11$ $0.76\pm 0.12$
+MB-SSd-4-1000-10 $0.13\pm 0.012$ $0.63\pm 0.1$ $4.7\pm 0.78$
+MB-SSd-4-1000-16 $0.067\pm 0.0088$ $0.39\pm 0.08$ $5.8\pm 1.2$
+MB-SSd-4-1000-22 $0.015\pm 0.0041$ $0.17\pm 0.064$ $11\pm 4.3$
+MB-SSd-4-1600-10 $1.7\pm 0.045$ $2.4\pm 0.21$ $1.4\pm 0.13$
+MB-SSd-4-1600-16 $0.52\pm 0.025$ $1.3\pm 0.16$ $2.6\pm 0.3$
+MB-SSd-4-1600-22 $0.39\pm 0.021$ $1.7\pm 0.17$ $4.3\pm 0.44$
+MB-SSd-4-2200-16 $0.77\pm 0.03$ $1.4\pm 0.16$ $1.8\pm 0.21$
+MB-SSd-4-2200-22 $1.6\pm 0.042$ $3.9\pm 0.26$ $2.5\pm 0.17$
+MB-SSd-4-2800-16 $1.1\pm 0.036$ $0.99\pm 0.14$ $0.92\pm 0.13$
+MB-SSd-4-2800-22 $1.9\pm 0.047$ $4.6\pm 0.3$ $2.4\pm 0.16$
+Table 6: Event yields for the squark pair production, $m_{\tilde{q}}=1700$ GeV, $m_{\tilde{\chi}_{1}^{0}}=500$ GeV with MC statistical uncertainties, $s\pm\Delta s$, normalized to the nominal cross section (8-fold degeneracy) at the approximate NNLO+NNLL accuracy, $\sigma=0.74$ fb.
+Bin CheckMATE ATLAS ATLAS/CheckMATE
+MB-GGd-4-1000-10 $4\pm 0.13$ $2.9\pm 0.49$ $0.74\pm 0.12$
+MB-GGd-4-1000-16 $1.4\pm 0.076$ $0.64\pm 0.21$ $0.46\pm 0.15$
+MB-GGd-4-1000-22 $18\pm 0.027 0.0$$097\pm 0.097$ $0.55\pm 0.55$
+MB-GGd-4-1600-10 $17\pm 0.26$ $17\pm 1.3$ $1\pm 0.076$
+MB-GGd-4-1600-16 $13\pm 0.23$ $4\pm 1.1 1.1$$1\pm 0.086$
+MB-GGd-4-1600-22 $5.1\pm 0.15$ $4.4\pm 0.63$ $0.85\pm 0.12$
+MB-GGd-4-2200-10 $11\pm 0.21$ $13\pm 1.1$ $1.2\pm 0.1$
+MB-GGd-4-2200-16 $14\pm 0.24$ $17\pm 1.2$ $1.2\pm 0.088$
+MB-GGd-4-2200-22 $12\pm 0.23$ $4\pm 1.1 1.1$$1\pm 0.086$
+MB-GGd-4-2800-10 $2.6\pm 0.11$ $2.4\pm 0.42$ $0.91\pm 0.16$
+MB-GGd-4-2800-16 $3.6\pm 0.12$ $3.6\pm 0.54$ $1\pm 0.15$
+MB-GGd-4-2800-22 $4.3\pm 0.13$ $5.3\pm 0.65$ $1.2\pm 0.15$
+MB-GGd-4-3400-10 $0.49\pm 0.046$ $0.56\pm 0.21$ $1.1\pm 0.42$
+MB-GGd-4-3400-16 $0.47\pm 0.046$ $0.3\pm 0.14$ $0.64\pm 0.3$
+MB-GGd-4-3400-22 $0.71\pm 0.056$ $1.5\pm 0.36$ $2.1\pm 0.51$
+MB-GGd-4-4000-10 $0.11\pm 0.023$ $0.24\pm 0.14$ $2.2\pm 1.3$
+MB-GGd-4-4000-16 $1\pm 0.022 0.11±0$$1\pm 0.1$ $0.98\pm 0.95$
+MB-GGd-4-4000-22 $0.21\pm 0.031$ $0.1\pm 0.1$ $0.48\pm 0.48$
+Table 7: Event yields for the gluino pair production, $m_{\tilde{g}}=1800$ GeV, $m_{\tilde{\chi}_{1}^{0}}=800$ GeV with MC statistical uncertainties, $s\pm\Delta s$, normalized to the nominal cross section at the approximate NNLO+NNLL accuracy, $\sigma=2.93$ fb.
+Bin CheckMATE ATLAS ATLAS/CheckMATE
+MB-C-2-1600-16 $2.2\pm 0.05$ $1.9\pm 0.18$ $0.89\pm 0.084$
+MB-C-2-1600-22 $9\pm 0.066 3.3$$3\pm 0.24$ $0.85\pm 0.062$
+MB-C-2-2200-16 $2.1\pm 0.049$ $2.3\pm 0.22$ $1.1\pm 0.1$
+MB-C-2-2200-22 $9.1\pm 0.1$ $9.8\pm 0.42$ $1.1\pm 0.047$
+MB-C-2-2800-16 $1.7\pm 0.044$ $1.5\pm 0.17$ $0.91\pm 0.098$
+MB-C-2-2800-22 $9.5\pm 0.1$ $9.5\pm 0.41$ $1\pm 0.044$
+MB-C-4-1600-16 $0.76\pm 0.029$ $0.71\pm 0.12$ $0.93\pm 0.16$
+MB-C-4-1600-22 $0.69\pm 0.028$ $0.63\pm 0.11$ $0.91\pm 0.16$
+MB-C-4-2200-16 $\pm 0.034 1.1$$\pm 0.14 1.1$$1\pm 0.14$
+MB-C-4-2200-22 $2.9\pm 0.058$ $2.5\pm 0.21$ $0.86\pm 0.07$
+MB-C-4-2800-16 $1\pm 0.035$ $0.76\pm 0.13$ $0.73\pm 0.12$
+MB-C-4-2800-22 $4.7\pm 0.074$ $4\pm 0.28$ $0.84\pm 0.06$
+MB-C-5-1600-16 $0.35\pm 0.02$ $0.28\pm 0.07$ $0.8\pm 0.2$
+MB-C-5-1600-22 $0.19\pm 0.015$ $0.24\pm 0.068$ $1.2\pm 0.35$
+MB-C-5-2200-16 $1\pm 0.034$ $0.61\pm 0.11$ $0.61\pm 0.11$
+MB-C-5-2200-22 $1.6\pm 0.043$ $1.2\pm 0.15$ $0.77\pm 0.094$
+MB-C-5-2800-16 $1.5\pm 0.043$ $0.75\pm 0.13$ $0.5\pm 0.085$
+MB-C-5-2800-22 $4.6\pm 0.074$ $3.3\pm 0.25$ $0.71\pm 0.055$
+Table 8: Event yields for the squark pair production, $m_{\tilde{q}}=1700$ GeV, $m_{\tilde{\chi}_{1}^{0}}=500$ GeV with MC statistical uncertainties, $s\pm\Delta s$, normalized to the nominal cross section (8-fold degeneracy) at the approximate NNLO+NNLL accuracy, $\sigma=0.74$ fb.